Use the following graph for a private closed economy to answer the next question.

$500 C+I 450 E 400 350 a 300 250 으 200 E 150 100 U 50 45° $50 100 150 200 250 300 350 400 450 500 550 Real Domestic Output ($B) 0 3 9

The multiplier for the above economy is:

A. 2

B. 3

C. 4

D. 5
Answer #1

The formula for calculating the Multiplier in a closed economy is=1/mps.

where MPS is the Marginal Propensity to Save.

  • If you know the MPC, you know the MPS, because MPC + MPS = 1.

mpc is marginal propensity to consume= ΔC / ΔY, where ΔC is change in consumption, and ΔY is change in income

\DeltaC=150 \DeltaY=300.

mpc=150/300=.5

mps=1-.5=.5

multiplier=1/.5=2